OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (e.mccormick):

Do you know the point slope formula?

OpenStudy (anonymous):

no

OpenStudy (e.mccormick):

\(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

so what would be your best answer

OpenStudy (e.mccormick):

Well, you are given the point \((2, -4)\) and slope \(\frac{3}{5}\). So the \((x_1,y_y)\) and m are given to you right there. Put them into that formula.

OpenStudy (e.mccormick):

Just be careful with the \(y_1\). The minus a negative becomes a positive.

OpenStudy (anonymous):

sorry still can't figure it out

OpenStudy (e.mccormick):

Well, what does \(y-y_1=m(x-x_1)\) become when you put your given numbers into it?

OpenStudy (e.mccormick):

\[\text{Given:}\\ (2,-4)=(x_1,y_1)\\ m=\frac{3}{5}\\ \therefore y-y_1=m(x-x_1) = \text{What?}\]Jut put things in place and that is a formula for your line. Done. Answered. Now, if you need to do something else, like get it into slope-intercept form, that may take a little more. But if all you need is a valid equation for the line, the answer is as simple as putting those three things into the formula and it is done!